What is the answer to:
7y-(6y+3)=12

Answer by checkley71(8403)

7Y-(6Y+3)=12
7Y-6Y-3=12
Y=12+3
Y=15
PROOF
7*15-(6*15+3)=12
105-90-3=12
105-93=12
12=12